OpenStudy (anonymous):

A toddler’s playground has two slides on either side. The slide on the left is situated at a 45° angle to the ground and is 52 centimeters away from the base of the playground. The slide on the right is positioned at a 30° angle to the ground and is 98 centimeters away from the playground’s base. What is the total length of both slides on this playground rounded to the nearest tenth of a centimeter?

OpenStudy (anonymous):

@ParthKohli Can you help her with this?

Parth (parthkohli):

Hmm can you make a diagram for me? I'm not good at understanding verbally :(

Parth (parthkohli):

Oh this is pretty easy. Do you know trigonometric ratios? If not I'll explain this in another way.

Parth (parthkohli):

I prefer trigonometric ratios btw.

OpenStudy (anonymous):

I don't. I need this question too.

OpenStudy (anonymous):

Would the answer happen to be 186.7 or is that completely off?

OpenStudy (anonymous):

you can just use cosine to find 2 lengths

OpenStudy (anonymous):

well i agree trig ratio is faster but cosine is easier

Parth (parthkohli):

Okay, we have to determine the hypotenuse in both triangles. In the slide towards the left, we'll use cos. \(\Large \color{Black}{\Rightarrow cos45 = {45 \over x}}\)

OpenStudy (anonymous):

Kk

Parth (parthkohli):

xsec45 = 45 We get sqrt2/1 as the sec45, so the hypotenuse would be sqrt2 * 52 = 52sqrt2

OpenStudy (anonymous):

Got that.

Parth (parthkohli):

Then the right slide. You use cos again.(cos of 30)

OpenStudy (anonymous):

98/cos(30)

OpenStudy (anonymous):

What do I add up now? @ParthKohli

Parth (parthkohli):

Oops my browser crashed

OpenStudy (anonymous):

I'm sorry. That always sucks.

OpenStudy (anonymous):

Could you show me how to get the answeR? I have to go to bed soon.

OpenStudy (anonymous):

This is my last question.

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow cos30 = {98 \over x} }\) Cos30 = sqrt3/2, so \(\Large \color{Black}{\Rightarrow {\sqrt3 \over 2} = {98 \over x} }\) \(\Large \color{Black}{\Rightarrow x\sqrt3 = 196 }\) \(\Large \color{Black}{\Rightarrow x = {196 \over \sqrt3} }\) \(\Large \color{Black}{\Rightarrow x = {196\sqrt{3} \over 3} }\)

Parth (parthkohli):

Well that's it

Parth (parthkohli):

Wait a second. That answer should be 196...

Parth (parthkohli):

196 + 52sqrt2

Parth (parthkohli):

I did a mistake there.

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow x\sqrt3 = 196}\) Yes, I was wrong

Parth (parthkohli):

The answer is 196 + 52sqrt2